### v3

# Field-free, Quasi-continuous Operation of Optical Nanofiber Interface with Two-dimensional Ferromagnetic Trap

Abstract
A soft ferromagnetic foil uniformizes Tesla-level magnetic fields generated by attached permanent magnets, producing a uniform and electronically tunable surface field on the opposite side. By arranging $n$ precisely fabricated rectangular foils, a nearly ideal magnetic quadrupole field with a substantial gradient can be created at center. This robust and tunable field configuration is useful for 2-dimensional magneto-optical trapping (2D-MOT) and magnetic guiding of cold atoms. In this work, by aligning an optical nanofiber (ONF) to the zero-field line of a 2-foil-based planar 2D-MOT, we demonstrate field-free operation of the quantum optical interface in a quasi-continuous manner, without switching off the magnetic field. Transient transmission spectroscopy is performed with a measurement repetition rate as high as 250 kHz. An anomalous line broadening is observed, which is not fully understood, but is partly explained by a small residual field along the zero-field line. Through additional field measurements and simulations, we clarify that this residual field can be eliminated in an $n$=4 assembly, resulting in an ultra-straight 2D trap to support efficient sub-Doppler cooling and uniform light-atom interaction over exceptionally long field-free distances $l$. With the strong field gradient to support atom guiding, the ferromagnetic device may also enable new quantum optical scenarios featuring interactions between co-guided atoms and photons.
I Introduction
Optical nanofibers (ONFs) have become integral components in the field of quantum optics [1, 2, 3, 4, 5, 6], primarily due to their ability to support highly efficient interactions between guided photons and surrounding near-field atoms. These interactions are sustained over exceptionally long distances, denoted as $l$, which are critical for realizing strong, collective light-atom coupling. However, achieving such precise interactions necessitates maintaining the ONF interaction zone free from any uncontrolled level-shifting fields, which can otherwise distort the atomic energy levels and broaden the coupling strengths. Additionally, keeping the ONF interface field-free is important for precision atomic spectroscopy at the nanoscale [7, 8], and for sub-Doppler cooling [9] which is an essential step to load a lattice [10, 11, 12, 13, 14]. However, laser cooling starts with magneto-optical trapping (MOT) [9]. For efficient spatial confinement of the atomic sample, MOT operation requires magnetic field gradients on the order of 10 G/cm. This leads to Gauss-level average fields along the ONF interface even for $l$ of merely a few millimeters. Switching off the MOT field to prevent such perturbations can be too slow in many experimental setups [15, 16, 10, 8, 2, 17], particularly since these measurements are often at single photon level and requires many repetitions for enough statistics.
A potential solution to meet the field-free requirements without switching off the field is to load the ONF interface with a two-dimensional MOT [18]. As in Fig. 1a, 2D-MOT operates with magneto-optical trapping force in the $x-y$ plane only, without requiring any field gradient along $z$. Cold atoms trapped in 2D-MOT is highly elongated, naturally serve to load the ONF interface. By precisely locating the ONF to the zero-field line, the magnetic field sensed by cold atoms in the nanoscale near field is practically zeroed, even for 2D-MOT operates with very high field gradients [19]. Importantly, since polarization gradient cooling is most efficient at zero field [20, 18], we expect efficient loading of atoms to the ONF optical lattice [10, 12, 13, 14]. However, all these 2D-MOT benefits require precise alignment of the zero-field line with the long ONF. While ONF is itself highly stiff [5] to support a straight line within e.g. a centimeter, a 2D-MOT optimized for free-space quantum optics [21] does not guarantee a straight enough zero-field line to match the straight ONF. In addition, to construct a 2D-MOT usually requires closely spaced rectangular coils carrying high currents [21], a setup that may be challenging to implement to ONF-based quantum optical platforms [22, 23, 24] where the ONF mounting structures necessarily constrains the size and complexity of the setup.
Figure 1: (a): Schematic of an ONF-2D-MOT interface with interaction length $l$. (b): Uniformization of magnetic field by a rectangular ${\bf\mu}-$metal foil. For the precisely machined foil, the magnetic field within the dashed line box on top can be highly uniform, being nearly invariant along $z$. The center of the $S_{2}$ surface ${\bf r}_{s}$ is marked with a red triangle marker. (c): The prism 2D-MOT using the $n$=2 foil structure in this work. Local laser beam polarizations defined along $+x,+y$ axes are marked. (d): 2D-MOT using $n$=4 foil structure. Local laser beam helicity are defined along $+x,+y$ axes in (a,c), and along a pair of $45^{\circ}$ axes in (d). In (c,d) the blue disk in the MOT center represents ONF. Its sub-micron diameter is exaggerated for clarity.
In this work, we draw inspiration from earlier developments in magnetically guided atom interferometry [19] and use ferromagnetic trap for field-free operation of ONF interfaces. As in Fig. 1(b), soft-ferromagnetic ${\bf\mu}-$metal foil poled by permanent magnets from one side can generate highly uniform surface field on the other [25, 26]. By properly arranging $n=2$ (Fig. 1c) or $n=4$ (Fig. 1d) foils, a zero-field line is formed near the symmetry center to support the 2D-MOT operation. This zero-field line is ultra-straight [19]. Its position can be finely adjusted via the electronic current surrounding the foils. By shifting this zero-field line to overlap with the straight ONF, the magnetic field is nullified microscopically to support field-free operation without switching off the magnetic field for the 2D-MOT. We experimentally demonstrate the key ingredients of our proposal with a compact two-foil ONF interface, featuring microsecond-level cooling-measurement cycle time. An anomalous line-broadening [10] in transient ONF spectroscopy is observed, which is not fully understood, but is partly attributed to a residual field gradient along the zero-field line direction $z$ (Fig. 2c) in the $n=2$ assembly. This $z$-gradient can be completely eliminated in the $n=4$ configuration (Fig. 2(a,b)), with which we discuss on the prospects for quasi-continuous, field-free operation of ONF-cold-atom interface with $l\geq$100 mm interaction distances, and to enable novel quantum optical research featuring strongly coupled photons with matterwave in their respective waveguides.
In the following our work is presented in three sections. First, in Sec. II, we outline the general principles of constructing an ultra-straight ferromagnetic trap for 2D cooling and trapping and for integrating the device with an optical nanofiber. In Sec. III, we present our experimental demonstration on field-free operation of the ONF interface, in particular, on a quasi-continuous transient absorption spectroscopy measurement with a repetition rate as high as 250 kHz. In Sec. IV we clarify practical aspects of the ultra-straight 2D field generation and propose an ONF interface with up to $l=10~{}$cm field-free interaction length. We summarize this work in Sec. V with an outlook into prospects of combining the ONF-interface with the ferromagnetically generated, nearly ideal 2D trapping potential.
II Ferromagnetic assembly for 2D cooling and trapping
II.1 Robust uniformization of magnetic field
The ${\bf\mu}-$metal is a nickel-iron-based soft-ferromagnetic alloy featuring high permeability ${\bf\mu}={\bf\mu}_{r}{\bf\mu}_{0}$, high saturation field $B_{\rm sat}$, and low remanence field $B_{\rm rem}$. The vacuum permeability is ${\bf\mu}_{0}=4\pi\times 10^{-7}~{}{\rm N/A^{2}}$. For example, the ${\bf\mu}-$metal material considered in this work is composed of 80% Nickel and 15% Iron. It has a saturation field of $B_{\rm sat}\approx 0.8$ T, a permeability of ${\bf\mu}_{r}\sim 10^{5}$ at $B\ll B_{\rm sat}$, and $B_{\rm rem}\sim$ 1 Gauss [26].
As in Fig. 1b, we consider rectangular ${\bf\mu}-$metal foil with width $W$, length $L$, and thickness $d\ll W,L$. The foil is magnetized primarily by permanent magnets, e.g. a Neodymium Iron Boron (NdFeB) block, from the $S_{1}$ side. To provide electronic tunability, the foil is in addition poled by current-carrying coils surrounding the foil. The NdFeB magnet with size $a\approx d$ is small. In absence of the ${\bf\mu}-$metal, its field at distance $W$ away is effectively a dipole field which varies substantially over any $O(W)$ distance. However, as highlighted by the dash-lined box in Fig. 1b, the field flux after being guided through the ${\bf\mu}-$metal becomes uniform near the distant surface $S_{2}$. Here, ${\bf r}_{h}={\bf r}_{s}+(x_{h},y_{h},z_{h})$ with $|x_{h}|,|y_{h}|=O(h)\ll W$ from the foil surface, and $|z_{h}|<L/4$ to avoid edge effects.
[…]
Here, we provide an intuitive explanation on the robust field uniformity, which is analogous to field uniformization by an isolated conductor in electrostatics [25]. Specifically, the ${\bf\mu}-$metal foil responses to the source field with a redistribution of effective magnetic charge. Similar to electrostatic shielding by a grounded conductor, the magnetic charge redistribution shields the inhomogeneous source field near $S_{1}$ from affecting the proximity of $S_{2}$. The shielding is therefore highly efficient for foil with large enough $W$, $L$, and $d$. Certainly, to keep the magnetic charge neutrality, there are extra charges in the ${\bf\mu}-$metal, which tend to distribute uniformly on the surface to minimize the field energy [25]. This surface charge picture becomes accurate near $S_{2}$ where $B\ll B_{\text{sat}}$ so that the ${\bf\mu}-$metal response is linear. In addition, with ${\bf\mu}_{r}\gg 1$ in the ${\bf\mu}-$metal, the ${\bf B}$ field in air (or vacuum) is perpendicular to the surface, according to the tangential ${\bf H}=\frac{1}{{\bf\mu}}{\bf B}$ continuity. As in electrostatics, the uniform surface charge leads to highly uniform ${\bf B}$ near the flat $S_{2}$ surface, insensitive to small variations of source fields or the ${\bf\mu}_{r}$ distribution. The uniformization is also effective against imperfect current-coil winding [19]. This effect is most pronounced for thin-wire winding at a sufficient distance from the $S_{2}$ surface, as illustrated in Fig. 2(c,d). While the shielding of source magnetic irregularities tends to degrade at large working distances, the shielding can be reinforced, if necessary, by constructing additional, ${\bf\mu}-$metal enclosures for the sources (Fig. 2d).
Finally, the field distribution is sensitive to the surface shape of the ${\bf\mu}-$metal itself, which can be machined with micrometer precision, over large scale. Mesoscopically, surface corrugations with a characteristic length scale $\xi$ of micrometers are typically present in these foils, but the associated fringe fields decay exponentially along $y$ [27, 28]. Therefore, to mitigate the effects of fringe field irregularities, 2D trapping should be positioned at a distance $y_{h}=h\gg\xi$.
Figure 2: Numerical simulation of the 2D-quadruple field [29, 30] for the 2-foil assembly, (a,b), according to Sec. III.1; and 4-foil assembly, (c,d), according to Sec. IV.1. See Appendix A for simulation details. The foil dimensions are not to scale. A.S.: Optional, additional shielding of source magnet variations for operating 2D-trap at large $h,h^{\prime}$ distances [31].
II.2 $n=2$ foil trap
As in Fig. 1c, Fig. 2(a,b) and Fig. 3, we consider arranging $n=2$ foils symmetrically in the $x-z$ plane with a gap distance of $\Delta y=2h$. With the opposite magnetization, the fields generated by the two foils cancel at ${\bf r}_{0}=(0,0,0)$ to form a magnetic zero. Near the center, the field ${\bf B}=B_{1}^{x}x{\bf e}_{x}+B_{1}^{y}y{\bf e}_{y}+B_{1}^{z}z{\bf e}_{z}$, with gradients $B_{1}^{x}:B_{1}^{y}:B_{1}^{z}=(1-\varepsilon):-1:\varepsilon$. Here $\varepsilon\sim(h/L)^{2}$ when $\varepsilon\ll 1$. The field zero is therefore a three-dimensional (3D) one. The residual $B_{1}^{z}$ gradient is associated with the fact that the field contributions from the two foils add up along $z$ for off-center locations. Nevertheless, with an operation distance $\varepsilon=h/L\ll 1$, the field distribution becomes highly elongated along $z$ to mimic a 2D field, as suggested by the dashed-line box in Fig. 2(c,d), with an approximate zero-field line near the center.
II.3 $n=4$ foil trap
As in Fig. 1d and Fig. 2(c,d), we now consider arranging $n=4$ foils symmetrically in the $x-y$ plane, with a gap distance of $\Delta y=2h$ and $\Delta x=2h^{\prime}$ (Fig. 2b). As in Fig. 2c, in contrast to the $n=2$ case, in the $n=4$ assembly the field variation along $z$ caused by each foil due to the finite $L$ are completely canceled at $x=y=0$. This results in a quadrupole field distribution, ${\bf B}=B_{1}y{\bf e}_{x}+B_{1}x{\bf e}_{y}$ at the center. Furthermore, with $h,h^{\prime}\ll W$, the gradient $B_{1}$ varies extremely slowly within e.g. $|z|<L/4$, as suggested by the dashed-line box in Fig. 2(c,d).
II.4 Electronic tuning
An important advantage of soft ferromagnets is that their magnetization is electronically tunable. As suggested by Fig. 1 and demonstrated with the Fig. 3f data, a slight change of currents surrounding the foils offsets the ${\bf\mu}-$metal magnetization. By adjusting the currents surrounding the selected foils in the $n-$foil assembly, the zero-field line can be shifted in a precise and repeatable manner. As an extension of previous ferromagnetic trapping techniques [18, 19], hybridizing the permanent magnetization with electronic tuning results in strong and adjustable quadruple field gradient $B_{1}$ at large distance $h$, without requiring high electronic currents. Of course, the disadvantage of our approach here is that one cannot switch off the permanent magnets. Fortunately, since the ONF-2D-MOT interface only samples the nanoscale interaction zone, one only needs to align the zero-field line to the ONF location in the $x-y$ plane, for the field-free operation of atomic physics and/or quantum optics experiments. The magnetic field does not need to be switched off globally.
II.5 2D-MOT-based ONF interface
Figure 3: (a): Schematic of the ONF-2D-MOT apparatus in this work. (b): The ONF-2D-MOT interface through the vacuum window. The ONF is visible through Rayleigh scattering of a 200 $\rm{\bf\mu}W$ heating light at 843 nm. (c): The 2D-MOT (marked with red arrow) and its three mirror images. (d): Schematic of the ONF-2D-MOT in planar geometry aided by two pairs of right-angle prisms. (e): Time-dependent fluorescence count through ONF and effective atom number $N_{\rm eff}$ during the MOT loading. The fit model for loading the density-limited MOT [20] (Appendix B), with $\rho_{0}=4\times 10^{10}/{\rm cm}^{3}$ final density near ONF, suggests a MOT lifetime of $\kappa^{-1}=0.4~{}{\rm sec}$. (f): Steady-state fluorescence count as a function of 2D-MOT position, adjusted by the $x$-coil and the foil-coil currents.
Both the quasi-2D field as in Fig. 2(a,b) and the 2D field as in Fig. 2(c,d) can be combined with properly polarized laser beams for magneto-optical trapping.
For the two-foil assembly, the easiest way to form the quasi-2D-MOT is to use a right-angle prism pair that retro-reflect a single beam (Fig. 1c). The confinement along the $z-$direction is provided by two additional beams, also with two prisms in this work (Fig. 3d). This planer arrangement can be conveniently integrated with an ONF mount [13] (Fig. 3(a-c)), as suggested by Fig. 1c. By adjusting the foil-coil current to shift the field along $y$, and an $x-$coil current to shift the field along $x$, the zero-field line is precisely adjusted in the $n=2$ assembly to overlap with ONF. Clearly, the two-foil configuration only supports a quasi-2D-MOT. Depending on the $\varepsilon=h/L$ ratio, the weak gradient along the ONF direction may limit the field-free ONF-interaction length $l$.
On the other hand, as illustrated with Fig. 2(c,d), a true 2D quadruple field can be formed with the $n=4$ foil assembly. Similar to Fig. 1c, one prefers a planer geometry for the 2D-MOT construction, which can then be integrated with a standard ONF-mount [13]. Here, with the field lines rotated by $45^{\circ}$ comparing to the $n=2$ case, a 2D mirror MOT [19] is likely preferred for optimal optical accesses. The ONF can be integrated on top of the mirror along $z$. The zero-field line can be shifted freely in 2D by adjusting, e.g., the currents surrounding the foil pair on the left of Fig. 2d.
III Experimental demonstration
We take advantage of two-foil compactness (Fig. 1c) to construct a planar ONF-2D-MOT interface, schematically illustrated in Fig. 3. We further demonstrate transient absorption spectroscopy measurement with a measurement repetition rate of $f_{\rm rep}=250~{}$kHz. An anomalous line-broadening is found which is partly explained by the residual $B_{1}^{z}$ gradient intrinsic to the two-foil assembly (Fig. 2(a,b)).
III.1 Experimental Setup
III.1.1 Two-foil vacuum assembly
The experimental setup is illustrated with Fig. 3a. A tapered nanofiber is fabricated with standard method  [32, 33, 34, 35]. The transmission of ONF is quite high, $T>85\%$. The fiber is epoxied to a titanium mount. The central $l_{0}\approx 5$ mm ONF section is centered to the mount, beneath which two pair of right-angle prisms are epoxied to the surface for the MOT-beam retro-reflections. The central pair of prisms are 101010~{}10mm high. The two side prisms are 666~{}6mm high. The ONF mounting assembly, carefully chosen to be non-magnetic, is connected to a CF35 stainless steel (SS-304) conflate flange. The ONF is optically feedthroughed to the vacuum with a Teflon interfacing technique [36]. The whole vacuum assembly is maintained to a $10^{-7}$ Pa base pressure. By opening a valve of a rubidium reservoir, a 87Rb partial pressure at $10^{-6}$ Pa level is achieved near ONF. To avoid Rb condensation, we gently heat the titanium assembly to $30^{\circ}$. Similarly, the ONF is heated by 200 $\rm{\bf\mu}W$ of 843843843~{}843nm light, (Fig. 4a) which is combined with the 780780780~{}780nm resonant probe beam with polarization optics and holographic gratings.
The two-foil assembly is constructed outside the vacuum with foil dimensions of $L=120~{}\text{mm}$ and $W=40~{}\text{mm}$. As shown in Fig. 5, each ${\bf\mu}-$metal foil is magnetized by a pair of cylindrical NdFeB magnets (25 mm diameter, 10 mm height) positioned 60 mm apart along the $z$ axis. An inter-foil gap of $2h=70~{}\text{mm}$ is used, which exceeds the ideal condition $h\ll W$ due to the vacuum tube diameter. The relatively thick foils, with $d=10~{}\text{mm}$, are advantageous for the 2D-MOT operation distance of $h=35~{}\text{mm}$ in this ex-vacuo design. To allow dynamic adjustment of the magnetic field, each ${\bf\mu}-$metal foil is wrapped with 23 turns of insulated copper wire with 1 mm diameter. The current passing through one foil is subsequently referred to as the “foil-coil” current (see Fig. 3d). Additionally, a coil with diameter $D=280~{}\text{mm}$ is mounted 30 mm away to provide a $B_{x}$ bias field, with its current denoted as the “x-coil” current.
We characterize the field distribution with a combination of numerical simulation (Fig. 2(a,b)) and Gauss-probe measurements (Fig. 5). As detailed in Appendix A, by adjusting in the simulation the the NdFeB strength, the numerical results can match quite well with the measurements. The surface field near $S_{2}$ is highly uniform and stable (Fig. 5). For the two-foil assembly, we find a quasi-2D magnetic field-zero line near the ONF location with field gradients of $B_{1}^{x}\approx 21~{}{\rm G/cm}$, $B_{1}^{y}\approx-24~{}{\rm G/cm}$, $B_{1}^{z}\approx 3~{}{\rm G/cm}$ along $x,y,z$ respectively.
III.1.2 Prism MOT
[…]
III.2 MOT loading
The MOT loading process is monitored by the ONF fluorescence. As illustrated with Fig. 3d setup, we direct an ONF output to a holographic grating to select the 780 nm fluorescence for single-photon counting (SPCM-AQRH, Excelitas). We switch off and back on the MOT beams with an acousto-optical modulator (AOM) to record the MOT-filling process. From the camera image, the final length of 2D-MOT along $z$ and in the $x-y$ plane are $l=4~{}$mm and $\Delta x=0.8~{}$mm respectively.
The fluorescence counting rate curve as shown in Fig. 3e includes a rapid initial growth which is followed by a slow increase further. The underlying mechanism is explained as following: Magneto-optical trap is known to have an irregular density distribution [18] associated with the quite unavoidable variations of intensity-imbalance between counter-propagating MOT beams. During the initial MOT loading, atoms prefer to be accumulated in intensity-balanced locations. Later, when there are plenty of atoms in these locations, multi-scattering starts to drive their expansions, leading to a large, “density-limited ” sample [20] with a uniform local density $\rho_{0}$. Here, the rapid initially growth of ONF fluorescence is associated with filling of intensity-balanced spots across ONF. The slow increase of ONF fluorescence later is associated with the multi-scattering-driven spreading to uniformly fill the ONF interaction zone, which effectively increase the interaction length $l$. We fit the fluorescence curve with a phenomenological model to include this multi-scattering-driven MOT expansion dynamics, detailed in Appendix B. The multi-scattering-limited density $\rho_{0}$ is derived from ONF absorption measurement. In particular, we find that the fully loaded atomic sample attenuates the ONF-guided resonant probe with an ${\rm OD}=2.2$ optical depth (see e.g. Fig. 4b), which suggests a uniform $\rho_{0}=4\times 10^{10}$/cm3 over the $l=4~{}$mm interaction length. In Fig. 3e we also estimate $N_{\rm eff}={\rm OD}(t)/{\rm od}_{1}$ as an effective number of atoms coupled to ONF. Here ${\rm OD}(t)$ is scaled with the fluorescence counting rate during the MOT loading. The ${\rm od}_{1}=0.09$ is single-atom resonant optical depth at a 100100100~{}100nm distance away from the $d=500~{}$nm ONF [2, 37].
III.3 Field control
We shift the 2D-MOT along $x$ and $y$ by adjusting the $x$-coil and foil-coil currents, demonstrating this field control by measuring the ONF-coupled steady-state MOT fluorescence, as shown in Fig. 3f.
We begin by optimizing the ONF fluorescence counting rate through adjustments of both the $x$-coil and foil-coil currents. With the foil-coil current then fixed at its optimal 3.5 A, we scan the $x$-coil current in small steps between 0.76 A and 1.5 A. This procedure shifts the 2D-MOT along $x$ by approximately 1 mm, as indicated by the mirror images observed on the camera (upper-most and lower-most MOT images in Fig. 3c). Since the 2D-MOT has a transverse size of about 0.8 mm, this shift is sufficient to toggle the ONF-coupled fluorescence off and back on, as shown by the red markers in Fig. 3f.
Next, with the $x$-coil current set to its optimal 1.16 A, we scan the foil-coil current between 1.2 A and 5.4 A. This adjustment shifts both the 2D-MOT and its second-order image transversely by approximately 1.8 mm on the camera (the two central images in Fig. 3c). The resulting variations in the ONF fluorescence counting rate are represented by the blue markers in Fig. 3f.
From the data in Fig. 3f, we estimate the foil-coil current tuning response to be approximately 1 G/A at the 2D-MOT location. Notably, the field is enhanced by a factor of $\eta\approx 5$ due to the magnetization of the ${\bf\mu}$-metal. Without the ${\bf\mu}$-metal, the same foil-coil would produce only about 0.2 Gauss per Amp at the same distance. This soft-magnetic field enhancement leads to the low power operation. In this setup, $\eta\approx 5$ is constrained by the moderate $d/W$ ratio and the resultant strong depolarization field [25]. In future work, a shorter working distance $h$ and a thinner foil with a smaller $d/W$ ratio could reduce demagnetization, potentially increasing $\eta$ to the hundred level. Regardless of the $\eta$ value, the soft-magnetization protects field uniformity against the foil-coil winding imperfection, as being explained in Sec. II.1.
III.4 High-rep-rate transient ONF spectroscopy
Figure 4: (a): Absorption spectroscopy setup and timing sequence. SPCM: Single photon counting module. Here $\tau_{\rm c}=30~{}$ns. $\tau_{\rm p}=200~{}$ns, 180 ns in (b) and (c,d) respectively. (b): Typical SPCM readout of the ONF-transmitted probe (red) and reference (blue). (c): Steady-state absorption spectroscopy and Lorentzian fit. (d): Transmission of a resonant probe ($\Delta=0$) from (b) with fit according to linear dispersion theory  [38, 39, 25, 40]. The dashed box suggests the range of data for the average to compile the $\Delta=0$ data point in (c).
As illustrated in Fig. 2(a,b), even with a working distance of $h=35$ mm, the field distribution in the two-foil configuration remains highly elongated along $z$. The 3 Gauss/cm gradient leads to 0.6 Gauss field strength deviation across the $l=4$ mm ONF interaction section aligned to the quasi-zero-field line. This level of field suppression should be sufficient for atomic spectroscopy at moderate precision, allowing measurements without globally switching off the field.
Leveraging this near-zero-field feature, we design a cooling-probe sequence for high-speed atomic spectroscopy, as depicted in Fig. 4a, simply by modulating the laser beams. The measurement repetition period is $T_{\rm rep}=4~{}{\bf\mu}$s. Within each period, the cooling and repumping beams, along with the 843 nm ONF heating beam, are all turned off and then back on within 0.8 ${\bf\mu}$s. As in Fig. 4b, accounting for finite AOM switching time, $T_{\rm p}\approx 600$ ns is open for absorption measurements. This probe window is short enough to prevent significant atomic motion, enabling atoms to be recaptured by a 3.2 ${\bf\mu}$s cooling pulse for the next probe cycle.
During the probe window, we generate a composite probe pulse using a fiber electro-optical modulator (fEOM)-based optical arbitrary waveform generator (OAWG) [41]. The pulse includes a 30 ns “calibration pulse” followed by a $\sim 200$ ns “interrogation pulse”. The calibration pulse is 400 MHz blue-detuned from the $F=2\rightarrow F^{\prime}=3$ hyperfine transition, leaving its ONF transmission unaffected by the presence of cold atoms. The interrogation pulse is scanned in detuning $\Delta$ from -20 MHz to 20 MHz. We keep the probe power at a low level of 5 pW to avoid saturating the atomic transition [42]. For each detuning $\Delta$, we integrate the single-photon counting module (SPCM) measurements over $T_{i}=10$ seconds. The resulting single-photon counts from all $f_{\rm rep}\times T_{i}=2.5\times 10^{6}$ measurements are histogrammed to yield the ONF-transmitted probe signal $I_{p}(t)$, shown in red in Fig. 4b. We note the MOT fluorescence is visible in $I_{p}(t)$ beyond the probe time window, as expected.
Additionally, a portion of the probe pulse is diverted to a reference SPCM without passing through the ONF, providing a “reference” signal $I_{r}(t)$. By comparing the 30 ns “calibration pulse” between the probe and reference signals, we normalize $I_{r}(t)$ (displayed in blue in Fig. 4b). With the normalization, the time-dependent probe transmission during the interrogation time is obtained as $T(t)=I_{p}(t)/I_{r}(t)$ in Fig. 4d.
III.5 Anomalous line broadening
With the transient transmission data $T(t)$ as those in Fig. 4d, we retrieve the near-field coupled atomic response in two ways. First, as suggested by the green dashed box, we take the $T(t)$ data between $175~{}{\rm ns}<t<180~{}{\rm ns}$ and calculate the average $\bar{T}$ at the steady state. We then evaluate the optical depth of the evanescently coupled cold atoms according to Beer-Lambert law [43], ${\rm OD}=-{\rm ln}(\bar{T})$, at each probe detuning $\Delta$. Typical absorption spectroscopy data are given in Fig. 4c together with Lorentzian fit. In the second method, we choose a particular $T(\Delta,t)$ curve as those in Fig. 4b, and apply linear dispersion theory [38, 39, 25, 40] to fit the Lorentzian atomic response [44]. In both cases, we arrive at a fitted linewidth of $\tilde{\Gamma}\sim 2\pi\times 9\sim 10~{}$MHz. This value is substantially larger than one might expect from the $\Gamma=2\pi\times 6.1~{}$MHz natural linewidth of the D2 transition [45], even after accounting for the ONF-enhanced emission and the surface interactions [11, 17, 8].
We checked a list of potential systematics that may lead to the anomalous line broadening. First, it is worth noting that unlike previous spectroscopic work where a prolonged measurement could affect the atomic trajectories [10, 17, 8], here, the atomic motion according to MOT dynamics is effectively frozen for our nanosecond probe. By increasing the probe power and repeating the measurements, we verified that our probe pulse is well below saturation in the near field [42]. To verify that cooling is effective between the nanosecond probes, the results were also repeated by increasing $T_{\rm rep}$ to 8 ${\bf\mu}$s and 16 ${\bf\mu}$s. Finally, the linewidth is not noticeably affected by long-range resonant dipole interaction [3], which is verified by repeating the observation with reduced atom numbers.
The observed linewidth broadening is partly explained by residual magnetic field. In particular, we find numerically [46, 37] that coherent coupling of different Zeeman sublevels by Gauss-level field, perpendicular to the quantization axis set by light, leads to notable line-broadening. However, by shifting the 2D-MOT center in the $x-y$ plane as in Fig. 3f, and by scanning a $z-$bias field, we verified that the $\tilde{\Gamma}=2\pi\times 9\sim 10~{}$MHz linewidth is indeed as small as we can get. We also note the 2D-MOT and the nanofiber are parallel to each other within a few degrees (Fig. 3c). With ONF aligned to the field-zero line, the $B_{1}^{z}\approx 3~{}$G/cm gradient should merely lead to $\delta B=0.6$ G deviation of field strength. The average field strength is therefore at 0.3 G level, much less than what is required to broaden the line even by just 1 MHz, about 1 G, according to numerical simulations [46, 37].
Therefore, even though the $\tilde{\Gamma}\approx 2\pi\times 9\sim 10$ MHz linewidth could be partly explained by the residual magnetic field in our setup, it remains a puzzle to be fully explained. A detailed account of our investigation on this topic will be given in a future publication [47].
IV Toward an ultra-straight 2D ferromagnetic trap
As in Fig. 2(a,b), the field zero in the two-foil assembly is a 3D one. The field variation along $z$ only vanishes in the $\varepsilon=h/L\ll 1$ limit. On the other hand, this field variation is absent in the four-foil version, as being described in Fig. 2(c,d). Of course, the assumption is that the ${\bf\mu}-$metal foils are perfectly magnetized and aligned. In this section we check practical aspects of this assumption, by measuring the field distribution and compare with numerical simulations.
IV.1 Field uniformity and stability
[…]
The field measurement is performed with a Gauss probe with 1 mm sensor size and 0.50.50.5~{}0.5G resolution. The Gauss probe is mounted on a 3D translation stage, centered to the $S_{2}$ surface of the foil to be measured, and is gently pressed against the surface to fix the distance. Then, as in Fig. 5a, by translating the Gauss probe along $z$, the surface field is recorded across the $L=120~{}$mm foil length. The measurements are repeated over days to verify the field stability. Example data are given in Fig. 5b with diamond, triangular, and disk markers. The marker size is chosen according to the measurement uncertainty and our confidence in positioning the Gauss probe.
We first note that all the $B_{y}(z)$ field profiles in Fig. 5b align well with the expectations from numerical simulations [29, 30] shown in Fig. 5c (Appendix A). Slight but repeatable deviations in $B_{y}$ are observed on a $\sim 10~{}\mathrm{mm}$ scale, with a magnitude close to the 0.5 G resolution limit. These deviations are likely caused by corrugations along the edges of the $d=10~{}\mathrm{mm}$ $S_{2}$ surface, introduced during the chamfering process in these prototypes. Such effects can be mitigated in the future with improved machining techniques.
The measurement results are repeatable. Taking the 1-Foil-C data as an example, which consists of three repeated measurements conducted over the span of a week, the data deviation is less than 1 G, showing excellent overlap in the plot. On the other hand, given that NdFeB blocks exhibit variations in surface quality and magnetic strength, the overall surface field strength naturally varies from foil to foil, as illustrated in the examples in Fig. 5b. These variations can be managed by hand-picking NdFeB blocks with consistent properties or by fine-tuning the field strength using additional magnets.
We verify that the uniform field profile shown in Fig. 5b is insensitive to the way the source magnets are attached on the $S_{1}$ surface. Field variations at the $1\%$ level only occur near $S_{2}$ when one of the two NdFeB blocks is displaced on the $S_{1}$ surface by as much as a few millimeters—a displacement much larger than the positioning precision achievable with a millimeter-scale ruler. As discussed in Sec. II.1, the combination of efficient source-field shielding and surface field energy minimization enforces a uniform field distribution near $S_{2}$. For foils with large enough $W$, $L$ and $d$, field irregularities induced by variations of small magnets on the distant surface are suppressed. Similar shielding effects are expected to smooth out the irregularity in the foil-coil current [19].
Given the highly uniform surface field in the finely machined 4-foil assembly, the field distribution above $S_{2}$ for $h,h^{\prime}\ll L$ are also expected to be uniform (Fig. 2(c,d)), as dictated by field continuity [31]. In our prototype assembly, accurately mapping the 3D field distribution is technically challenging. Nevertheless, using the same Gauss probe, we measured the field both above the foil surfaces and near the zero-field region of the 4-foil assembly. These measurements align with simulations, albeit with reduced fractional resolution and increased positional uncertainties.
Finally, as in Sec. II.1, we emphasize the choice of large enough working distances $h$ for finely machined ${\bf\mu}-$metal foils, to avoid fringe fields associated with micrometer-scale ${\bf\mu}-$metal surface corrugations [27, 28].
IV.2 Alignment considerations
The construction of the 2D quadrupole field shown in Fig. 2(c,d) relies on precise alignment of the four uniformly magnetized ${\bf\mu}$-metal foils. In practice, field deviations can arise from misalignment or positional errors. However, because the ${\bf\mu}$-metal foils can be accurately machined and securely mounted, we anticipate small alignment or positional errors in the foil assembly will be manageable to induce negligible field deviations.
[…]
Next, we consider the effect of a positional error $\delta z$ along the $z$-axis by one of the foils. The resulting field perturbation can be interpreted as a shift in surface magnetic charge [25]. The unbalanced magnetic charge at the edges lead to a residual field at the center, which, for sufficiently large foil length $L$, becomes weak and smooth. Therefore, for precisely positioned foils with $\delta z/L$ at e.g. $10^{-4}$ level, and in addition $L\gg h$, this type of perturbation can be suppressed to milli-Gauss level while maintaining the $B_{1}\sim 100~{}$G/cm gradient.
On the other hand, for any one of the foils, small positional errors along the $x$ and $y$ directions shifts the zero-field line itself, and can be compensated by fine-tuning the magnetization through adjustments to the foil-coil current.
IV.3 Field-free ONF interface with $l\sim 100~{}$mm
So far, we have shown that it is practical to construct a 2D ferromagnetic trap as those in Fig. 2(c,d) with high precision. Enhanced by permanent magnets, a field gradient of $B_{1}=100~{}$G/cm can be quite easily achieved at $h=20~{}$mm distance for a four-foil assembly with e.g. $L=200$ mm. The center zero-field line is long enough to support ONF-2D-MOT system with $l\sim 100~{}$mm, with residual field suppressed to the milli-Gauss level. As such, the evanescent optical lattice [10, 11, 12, 13, 14] can be loaded from the high-gradient 2D-MOT without switching off the trapping field.
For a uniformly filled lattice, the collective interaction strength between guided photons with the lattice atoms increases linearly with $l$. With the $l\approx 100~{}$mm interaction length, the interaction strength would be quite unprecedented, which can be crucial for uncovering new regime in waveguide QED [43, 2]. Importantly, the quantum optical measurements can be performed in a quasi-continuous manner, as those demonstrated in this work (Fig. 4), leading to high data rate. The efficient laser-cooling in between measurements should also help to mitigate ONF vibration induced heating [48].
We finally note that while the magnetization and alignment can be optimized to improve the field-free operation of the ONF-interface, practically the level of field suppression along ONF is still likely limited by engineering imperfections. To this end, we highlight previous successful experience on ferromagnetic trap with guided matterwave interferometry [19]. We also note that for achieving improved interaction uniformity, the inhomogeneous broadening due to weak residual fields can be managed by spin echo and composite pulse techniques [49, 37].
V Summary and outlook
Integrating nanophotonics with cold atoms represents a promising route toward quantum information processing via nonlinear quantum optics [50]. A major technical challenge in this endeavor is the precise positioning of cold atoms at the nanophotonic interface. Among the most successful approaches to address this challenge is the ONF-cold atom interface, which has seen significant progress over the last decade [16, 10, 8, 2, 3, 17, 6]. In these works, sub-Doppler cooling is critical for efficient loading of the near-field lattice, while a uniform atomic spectroscopic response is essential to support long-range resonant dipole interactions. But the associated field-free requirement is at odds with the necessity of a strong field gradient for magneto-optical trapping. Managing these conflicting requirements becomes increasingly difficult at large $l$, the ONF interaction distance, which is preferred to be as long as possible. But switching off the field to accommodate these constraints severely limits the duty cycle of single-photon-level measurements.
In this work, we have proposed to resolve the contradictory field requirement at the cold-atom-ONF platform using 2D ferromagnetic trap [18, 19]. We highlight the unique opportunity of aligning the zero-field line in 2D-MOT with the stiff ONF, for efficient loading and field-free operation of the nanophotonic interface. We show that precisely fabricated ${\bf\mu}-$metal foils robustly uniformize source field from permanent magnets or current-carrying wires, for generating ultra-straight 2D field to match a long and straight ONF [5].
Experimentally, we integrate a two-foil ferromagnetic assembly with ONF using a planar 2D-MOT (Fig. 3). The field uniformity is confirmed with Gauss probe measurement on the foil surface (Fig. 5b). The 2D nature of the central quadrupole field (Fig. 2(a,b)) is verified by the robust operation of the highly elongated 2D MOT (Fig. 3e). Freed from the slow field-switching constraints, we perform atomic spectroscopy at a repetition rate as high as $f_{\rm rep}=250~{}$kHz, utilizing nanosecond probe pulses as shown in Fig. 4. We note that the quasi-continuous spectroscopic measurements minimally perturb the atomic trajectories [10]. In the ONF near field, the atomic density distribution is preserved as that during MOT operation, a situation not been carefully looked upon with precision spectroscopy previously, to our best knowledge. Therefore, the anomalous line-broadening effect in this unique system, observed in Sec. III.5, warrants further investigations [47].
Comparing with traditional ONF-interfaces, the ferromagnetic, planar 2D-MOT interface demonstrated in this work (Fig. 3a) is substantially simpler and more compact. The $\sim$20 G/cm gradient with sufficient tunability (Fig. 3f) is achieved at a notable long working distance of $h=35~{}$mm, facilitating the ex-vacuo operation. Based on field measurements and simulations, we provide arguments in Sec. IV that ONF interface with $l\approx 100~{}$mm can be constructed using a four-foil ferromagnetic trap. The unprecedentedly high light-atom coupling strength may help to unlock new research opportunities in waveguide QED [43, 2]. The quantum optical measurements can be performed in a quasi-continuous manner, as those demonstrated in this work. The efficient laser-cooling in between measurements should also help to mitigate ONF vibration induced heating [48].
Separately, with reduced $h$, the field gradient in these ferromagnetic traps can easily reach $B_{1}=100$ G/cm, strong enough for tightly confining laser-cooled atoms in their low-field-seeking states with magnetic dipole forces [19]. This may open up new quantum optical scenarios featuring waveguided photons interacting with magnetically guided atoms. To this end, we note related progresses are attempted with optical guiding inside hollow-core fibers [51, 52], where the field inhomogenuity poses the key challenge. Here, the ultra-smooth confining potential offered by the ferromagnetic trap would help to support matterwave coherence over a long distance comparable to that for the optical coherence in the ONF-based quantum optical platform [2, 3].
VI Acknowledgments
We are grateful to Professor Luis Orozco and Professor Wei Fang for valuable discussions. This project is supported by National Key Research Program of China under Grant No. 2022YFA1404204, from Natural Science Foundation of China under Grant No. 12074083, 12274272, from Natural Science Foundation of Shanghai under Grant No. 23dz2260100, and from the Shanghai Science and Technology Innovation Action Plan under the Grant No. 24LZ1400300.
Appendix A Magnetic field simulation
[…]
In the simulation, the ${\rm{\bf\mu}}$-metal parameters are chosen as ${\bf\mu}_{r}=10^{5}$, $B_{\rm sat}=0.8~{}\mathrm{T}$, and $B_{\rm rem}=0$ [26]. The surface field strength of the NdFeB block is set to $B_{0}=0.9~{}\mathrm{T}$, an empirical value selected to align the central field gradient of the assembly between measurements and simulations. The structural geometry corresponds to the experimental setups shown in Fig. 2 and Fig. 5.
The accuracy of the field simulation relies on the granularity of magnet segmentation and the numerical relaxation tolerance. We adjusted these parameters to ensure that neither finer subdivisions nor tighter relaxation tolerances significantly affected the final results. The relaxation tolerance is set to ${\rm prec}=10^{-4}$ T (1 G). For the simulations in Fig. 2 and Fig. 5c, each NdFeB block is divided into $10\times 10\times 10$ subunits. Each ${\rm{\bf\mu}}$-metal foil is divided into $n_{z}=160$ and $n_{y}=50$ segments along the $z$ (“$L$ direction”) and $y$ (“$W$ direction”) axes, respectively. Additional simulations confirmed that the central field profile $B_{y}(z)$, as shown in Fig. 5c, is insensitive to segmentation along the $x$-axis, allowing us to set $n_{x}=1$ to conserve computational resources. This approach neglects the variation of $B_{y}$ across the foil thickness ($x$), which is minimized at the center ($x=0$) and peaks at the edges. To account for this approximation, an overall scaling factor of 0.8 is applied to the field simulation in Fig. 5c, evaluated at $h=0.5~{}\mathrm{mm}$ from $S_{2}$, to match experimental measurements shown in Fig. 5b. This correction is necessary only in the vicinity of $S_{2}$. For field distributions where $h\gg d$, the simulation remains accurate, and we fix the NdFeB block surface field at $B_{0}=0.9~{}\mathrm{T}$ for all simulations.
Appendix B A multi-scattering model for loading the 2D-MOT-ONF interface
The process of loading a magneto-optical trap is characterized by the time-dependent atom number and the characteristic size of the MOT, $N(t)$ and $\sigma(t)$ respectively. For a single-chamber MOT where atoms are directly captured from the background vapor, as in this work, the $N(t)$ dynamics follows quite accurately the differential equation $\dot{N}=R-\kappa N-\beta N^{2}/\sigma^{3}$. Here the loading rate $R$ depends on the laser and magnetic field parameters, as well as the background rubidium partial pressure $P_{\rm Rb}$. Instead, the single-body loss rate $\kappa$ is almost solely decided by the total pressure $P$. The two-body loss coefficient $\beta$ describes the light-assisted collisional losses [53]. For the setup in this work (Fig. 3) with $P\approx 10^{-6}~{}$ascal, the $\kappa\sim 10~{}{\rm sec}^{-1}$ at moderate $R$ limits the number of atoms in the MOT and correspondingly the two-body loss. We therefore expect
$N(t)=N_{0}(1-e^{-\kappa t}).$ (1)
As being explained in Sec. III.2, for small enough $N(t)$ initially, the captured atoms tend to fill the equilibrium locations where radiation pressure is balanced for the nearly empty MOT. The size $\sigma_{0}$ for this temperature-limited MOT [20] is decided by the MOT restoring force and the equilibrium MOT temperature. Then, as $N(t)$ increases, the atomic sample becomes optically dense to substantially attenuate the MOT beams. This collective modification of radiation pressure is combined with other multi-scattering effects to limit the atomic density at the MOT center, by expelling additional atoms outward. Consequently, the density-limited MOT [20] accommodates more atoms by expanding its size $\sigma(t)$ with $N(t)$. For a moderate-sized MOT as in this work, we expect one-to-one mapping of $\sigma(t)$ and $N(t)$. That is, there is certain functional form
$\sigma(t)=\tilde{\sigma}(N(t),\rho_{0}),$ (2)
to decide the MOT size according to the atom number. The limiting density $\rho_{0}$ is largely decided by the intensity and detuning of the MOT beams.
We expect the Eqs. (1) (2) analysis to be valid for both 2D and 3D MOTs. Here, we assume the 2D-MOT size is isotropic in the $x-y$ plane, with $\sigma$ according to Eq. (2) and is $z-$independent. Furthermore, We assume the 2D-MOT length, $l=4~{}$mm, is time-independent. This is supported by the camera video (Fig. 3c) during the MOT loading. Physically, the stationary length is associated with the fact that the restoring force along $z$ in the (quasi)2D-MOT is too weak to support any interplay with the multi-scattering force.
The 2D-MOT is closely aligned to the zero-field line and ONF. But the center is corrugated, due to the MOT beam imbalance [18, 21]. For the 2D-MOT section that is $r_{\perp}$ away from ONF, the atom density at ONF is
$n(r_{\perp})=\frac{N}{\pi\sigma^{2}l}e^{-(r_{\perp})^{2}/\sigma^{2}}.$ (3)
Therefore, with the $r_{\perp}$ distance following certain $g(r_{\perp})$ distribution, the ONF-coupled MOT fluorescence counting rate scales as
$\gamma_{\rm SPCM}\propto\int n(r_{\perp})g(r_{\perp}){\rm d}r_{\perp}.$ (4)
To facilitate the numerical fit for Fig. 3c, in Eq. (2) we phenomenologically set $\tilde{\sigma}(N,\rho_{0})={\rm max}(\sigma_{0},\sigma_{s})$, with $\sigma_{s}=\sqrt{N/(l\rho_{0})}$ for the density-limited 2D-MOT. The density limit $\rho_{0}$ is fixed quite accurately first, by comparing the ONF-coupled resonant optical depth for the same MOT sample (Fig. 4) with the theoretical estimation based on optical scattering in the near field [46, 37], leading to $\rho_{0}=4\times 10^{10}/{\rm cm}^{3}$. By this step, we assume that after the MOT is fully loaded, $\rho_{0}$ is approximately reached in the near field of ONF.
We simply assume a uniform $g(r_{\perp})$ between $0\leq r_{\perp}\leq b$. With $b$ as auxiliary parameters, we fit the Fig. 3c data with Eq. (4) to retrieve $\kappa=(0.4~{}{\rm sec})^{-1}$. The fit suggests $\sigma_{0}\approx 0.6b$ for $t<0.1~{}$s, and then gradually expand to $\sigma(t)\approx b$ within one second. The ONF fluorescence is a local observable. Consequently, the atom number $N_{0}$ retrieved from the measurement depends on the $b$ parameter we choose in the model. For example, $N_{0}=3\times 10^{6}$ is reached by setting $b=100~{}{\bf\mu}$m.